shasha12:

A? What is the best approximation for the perimeter of a semicircle with a diameter of 64 meters? Use 3.14 to approximate pi. 100.48 m 164.48 m 200.96 m 264.96 m

Pixel:

pi x r + d

Pixel:

d = 64 r = d/2 r = 64/2

Pixel:

r = 32

Pixel:

3.14 x 32

Pixel:

100.48

Pixel:

+ d

Pixel:

100.48 + 64 = 164.48

Pixel:

B